Pompeii Pizza Club owns three identical restaurants popular for their specialty pizzas. Each restaurant has a debt-equity ratio of 30 percent and makes interest payments of $43,000 at the end of each year. The cost of the firm’s levered equity is 16 percent. Each store estimates that annual sales will be $1.29 million; annual cost of goods sold will be $750,000; and annual general and administrative costs will be $410,000. These cash flows are expected to remain the same forever. The corporate tax rate is 21 percent. a. Use the flow to equity approach to determine the value of the company’s equity. (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89) b. What is the total value of the company? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89)
Answer #1

1. value of company equity

particulars amount $

sales 1,290,000

cost of goods sold 750,000

general & admin. cost 410,000

interest expense 43,000

earning before tax 87,000

tax 18,270

net income 68,730

value of company's equity = 68,730/0.16

value of company's equity = $ 429,562.5

2. total value of company

debt equity ratio = 0.30

debt/equity = 0.30

debt / 429,562.5 = 0.30

debt = 429,562.5 * 0.30 = 128,868.75 $

value of company = debt + equity

value of company = 128,868.75 + 429,562.5

= $ 558,431.25